OpenStudy (anonymous):

Solve square root of the quantity x minus three is greater than or equal to seven. x ≥ 52 x ≤ 52 3 ≤ x ≤ 52 x ≥ 3

OpenStudy (anonymous):

\[\sqrt{x - 3}\ge 7\] Squaring, we get: x - 3 >= 49 Adding 3 to each side gives x >= 52

OpenStudy (anonymous):

Thank You